# Hyperfine spectroscopy of optical-cycling transitions in singly ionized thulium

### v1

@@ Abstract
We present a spectroscopic investigation of $^{169}\text{Tm}^{+}$ that provides two key foundations for its use as a platform for advanced quantum applications. First, we establish the complete spectroscopic road map for optical cycling (including laser cooling) by performing high-resolution spectroscopy on $^{169}\text{Tm}^{+}$ ions in an ion trap. We characterize the primary $313$ nm and complementary $448/453$ nm cycling transitions, identify the essential near-infrared repumping frequencies, and determine the magnetic-dipole hyperfine $A$ constants for all relevant levels. Second, we report detailed characterization of a metastable state as a candidate for hosting a robust qubit, performing lifetime measurements and Zeeman-resolved microwave hyperfine spectroscopy with $\mathrm{kHz}$ precision.
I Introduction
@@
Despite the maturity of neutral-atom quantum platforms across the periodic table [McClelland.2016], including alkalies [Isenhower.2010, Jau.2016], alkaline-earth-like atoms [Madjarov.2020, Ma.2022], rare-earth metals [McClelland.2006, Sukachev.2010, Lu.2011, Dutta.2014], metastable noble gases [Vassen.2012], and group III elements [Yu.2022], laser cooling and control in trapped ions have been mostly focused on alkaline-earth and alkaline-earth-like ions with an $ns\ {}^{2}\mathrm{S}_{1/2}$ ground state [Bruzewicz.2019]. While the level structure of such elements offers relatively convenient schemes for laser cooling, state preparation and measurement (SPAM), they lack a large number of accessible internal states and multilevel schemes. In contrast, additional long-lived states in non-alkali-like ions could enable architectures that partition roles across distinct qubits (e.g., processing vs. memory), as proposed in the optical, metastable, and ground-state (omgomg) qubit protocol [Allcock.2021, Vizvary.2024]. While the feasibility of laser cooling of open-$4f$-shell rare-earth ions has been discussed before [Lepers.2016], no real experimental efforts for laser cooling have been demonstrated. Here we present such efforts in $^{169}\text{Tm}^{+}$.
@@
Thulium is a rare-earth lanthanide metal with a single stable isotope, $^{169}\text{Tm}$. Neutral Tm has recently been laser cooled [Sukachev.2014, Provorchenko.2023] and used for atomic clocks [Sukachev.2016, Golovizin.2019], Bose–Einstein condensates [Khlebnikov.2019, Davletov.2020, Tsyganok.2023], and as a quantum-computing platform [Mishin.2025]. $^{169}\text{Tm}^{+}$ in its ground state has a single hole in the $4f$ shell and one electron in the $6s$ shell, generating the $4f^{13}6s\,{}^{3}\mathrm{F}_{4}^{\mathrm{o}}$ state, or $(\nicefrac{{7}}{{2}},\nicefrac{{1}}{{2}})^{\mathrm{o}}_{4}$ in the $J_{1}j$-coupling basis [NIST]. This configuration yields manifolds of relatively low-lying states with total electronic angular momentum spanning from $J=1$ to $J=9$. Such large-$J$ manifolds are attractive for absorption–emission code schemes that require at least $J\geq\nicefrac{{9}}{{2}}$[Jain.2024]. At the same time, the nuclear spin $I=\nicefrac{{1}}{{2}}$ ensures that each fine-structure level splits into only two hyperfine-structure (HFS) states, simplifying laser cooling and qubit initialization by reducing the number of frequency tones required for closed population transfer.
For $I=\nicefrac{{1}}{{2}}$ the HFS shift $\Delta\nu_{F}$ of a level with quantum numbers $(J,F)$ relative to its center of gravity is [Schwartz.1955]
$\Delta\nu_{F}=\frac{A}{2}\!\left[F(F+1)-J(J+1)-I(I+1)\right],$ (1)
where $\nu_{J}^{J^{\prime}}$ is the centroid transition frequency. Because the target manifolds are populated optically, the measured spectra also reflect population dynamics relevant to closing the cycles. Here we determine the HFS $A$ constants of excited levels in $^{169}\text{Tm}^{+}$ that are not directly accessible from the ground state via optical dipole transitions; these levels are populated via resonant optical population transfer within two candidate laser-cooling cycles.
II Spectroscopy methods
Of the spectroscopic data on $^{169}\text{Tm}^{+}$ that has been collected experimentally [Martin.1978, Mansour.1989, Wang.2022, Kebapci.2024, DenHartog.2024], only three transitions relevant to our proposed cooling schemes (described in Sec. IV) have been reported in the literature: $448.25$ nm, $453.07$ nm, and $313.22$ nm, with no data on the transitions corresponding to possible repumping transitions. Guided by theoretical predictions for possible decay channels [Quinet.1999, Quinet.DREAM, Palmeri.2025], we measured the transitions’ wavelengths and hyperfine structure of the relevant levels.
We employed three different experimental setups to determine the HFS of the levels involved in the proposed cycling schemes. Our most precise measurements were obtained from laser-induced fluorescence on a trapped ion cloud, while preliminary spectra were recorded using a hollow-cathode lamp and a small vacuum chamber with an ablation target. The latter two methods can be implemented quickly and cost-effectively compared with an ion-trap apparatus and are therefore well-suited for evaluating the feasibility of working with a new ion species before committing to modifications of an established trap setup.
II.1 Absorption spectroscopy in a hollow-cathode lamp
II.2 Fluorescence spectroscopy in an ablation cell
II.3 Fluorescence spectroscopy in a linear quadrupole trap
Figure 1: Image of an illuminated cloud of $\text{Tm}^{+}$ ions, taken with an EMCCD camera. Both the $448$- and $453$-nm lasers cross the trap center twice, forming an X shape. The image was post-processed with a median and a Gaussian filter to reduce image noise and improve contrast.
@@ II.3 Fluorescence spectroscopy in a linear quadrupole trap
Figure 2: Measured EE1 transitions and hyperfine splittings in $^{169}\text{Tm}^{+}$. Levels with the same $J$ value are aligned vertically, with $F$ values labeling the hyperfine sublevels. Level energies are indicated in $\text{cm}^{-1}$ and the measured hyperfine splittings are given in GHz. The level spacings are not to scale, and levels irrelevant to this work are omitted. The vacuum wavelength and corresponding frequency are shown next to one of the two strongest hyperfine components ($\Delta F=\Delta J$) of each transition. Gray levels and gray dash-dotted lines indicate levels and expected transitions, respectively, that were searched for but not observed in this work.
Most of our spectroscopy measurements were performed in a linear quadrupole trap originally designed for experiments with trapped $\text{Ba}^{+}$ ions [Putnam.2024, Putnam.2024.thesis]. The trap consists of four parallel, segmented rods arranged symmetrically around the trap axis at a distance of $1$ cm. Each rod is divided into three electrodes. A radio-frequency (RF) voltage of 0.50.511 kV at $0.939$ MHz is applied to a pair of opposing center segments to provide radial confinement. The fringe field of the RF potential also contributes to weak axial confinement; additional static voltages can be applied to the outer segments to displace the ion cloud or increase the axial trapping potential. The trap is situated inside a sealed chamber and can operate under ultrahigh vacuum conditions or with buffer gas.
The trapped ions are illuminated by multiple laser beams intersecting at the trap center, with up to two passes per beam. Each beam may contain several frequency components. The main cooling configuration for the $450$-nm colling cycle comprises two tones at $448$ nm and two at $453$ nm, crossing the trap center twice in an X-shaped geometry, as illustrated in FIG. 1. For the $313$-nm system, two counter-propagating beams each intersect the trap center twice in a “plus” geometry; the second beam is frequency-shifted by +722+722 MHz relative to the first using acousto-optic modulators (AOMs). The AOM shift ensures that both HFS levels are addressed. One infrared (IR) beam is counter-propagating the $450$-nm beam, and two additional IR beams intersect the trap center once.
Figure 3: Hyperfine-structure (HFS) spectra of $^{169}\text{Tm}^{+}$: (a,b) recorded using absorption spectroscopy with a hollow-cathode lamp, and (c–i) using fluorescence spectroscopy with an ionic cloud in a linear quadrupole trap. (c–f) A fresh ion cloud was produced for each data point; the laser light was generated using external-cavity diode lasers (ECDLs) and the laser frequencies were set to fixed values for each point. (g–i) Each spectrum was recorded with a single cloud of ions using the continuous-scan function of the optical parametric oscillator (OPO) laser.
All fluorescence spectra were recorded by collecting photons from the main cooling transitions at $448$, $453$, or $313$ nm. Potential repumping transitions were identified by monitoring their effect on this steady-state fluorescence. This technique works because the cooling lasers continuously populate intermediate “dark” states. When an additional laser is scanned into resonance with a repumping transition, it depopulates one of these metastable states and returns ions to the main cooling cycle. This results in a measurable increase in the steady-state fluorescence from the cooling transitions.
II.4 Laser systems
The wide range of wavelengths employed in this work requires different types of lasers. For infrared transitions from $1056$ to $1307$ nm, we used an optical parametric oscillator (OPO, Hübner C-wave GTR). For transitions between $749$ and $944$ nm, we used in-house built external-cavity diode lasers (ECDLs) or a Ti:sapphire laser (M-squared SolsTiS). Additional ECDLs were used to generate $448$ and $453$-nm light. The $313$-nm light was generated with a frequency-quadrupled diode laser (Toptica TA-FHG). All laser frequencies were measured with a wavelength meter (Angstrom/HighFinesse WS-8).
III Results
@@
The energy levels and EE1 transitions investigated and relevant for laser cooling and State Preparation And Measurement (SPAM) are depicted in FIG. 2. In this diagram, solid arrows indicate transitions observed experimentally, while dashed arrows denote potential decay channels predicted by theoretical calculations but not observed in our spectroscopic measurements. This level diagram is non-exhaustive; for clarity, we omit some intermediate levels and those with $J<3$ and $J>6$. For each wavelength, the measured frequency of one of the two strongest hyperfine components ($\Delta F=\Delta J$) is indicated together with the corresponding inferred level splittings.
@@
From the recorded fluorescence spectra, depicted in FIG. 3, the hyperfine-structure (HFS) $A$ constants of the excited states relevant to the proposed cooling schemes were determined. The extracted values are listed in Table 1. Each spectrum was fitted with a sum-of-Gaussians model, using the HFS $A$ parameters of the lower and upper states as fit variables. Due to the large total electronic angular momentum, transitions where $\Delta F\neq\Delta J$ are strongly suppressed. As a result, in several cases, only two spectral features were available to extract a single unknown $A$ constant, with the other fixed to a known reference value.
@@
Table 1: Hyperfine-structure (HFS) constants in $^{169}\text{Tm}^{+}$. The doublet splittings are readily calculated using $(J+\nicefrac{{1}}{{2}})A$. The HFS constants of the ground, $236.95$, and 31 926.8231\,926.82-cm-1 states were kindly provided by Ref. [Bodnar.2023]. The precise value of the 12 457.2912\,457.29-cm-1 state was determined using microwave spectroscopy.
@@
The specified uncertainties have been determined from the least-square-fit uncertainties, adjusted for a reduced $\chi^{2}$ of one, as well as the statistics of repeated measurements and the frequency readout. A statistical $5$ MHz frequency-readout uncertainty was added in quadrature (geometric sum) to the uncertainties of the determined resonance peak positions of spectra generated with the ECDLs, before calculating the weighted average over all measurements. The maximum between the standard errors of the unweighted and weighted average was then taken as the total statistical uncertainty. In cases where the $A$ constant was predominantly extracted from continuous scans of our OPO laser, a $2$ MHz systematic frequency-calibration uncertainty was added linearly to the statistical uncertainty.

### v2

@@ Abstract
We present a spectroscopic investigation of $^{169}\text{Tm}^{+}$ that provides two key foundations for its use as a platform for advanced quantum applications. First, we establish the complete spectroscopic road map for optical cycling (including laser cooling) by performing high-resolution spectroscopy on $^{169}\text{Tm}^{+}$ ions in an ion trap. We characterize the primary $313$ nm and complementary $448/453$ nm cycling transitions, identify the essential near-infrared repumping frequencies, and determine the magnetic-dipole hyperfine $A$ constants for all relevant levels. Second, we report a detailed characterization of a metastable state as a candidate for hosting a robust qubit, performing lifetime measurements and Zeeman-resolved microwave hyperfine spectroscopy with $\mathrm{kHz}$ precision.
keywords:
Thulium-169 ion spectroscopy, Hyperfine structure, Optical cycling transitions, Laser cooling of ions, Metastable qubit state, Ion trap experiments
Introduction
@@
Despite the maturity of neutral-atom quantum platforms across the periodic table [25], including alkalis [12, 14], alkaline-earth-like atoms [21, 20], rare-earth metals [24, 35, 19, 8], metastable noble gases [40], and group III elements [43], laser cooling and control in trapped ions have been mostly focused on alkaline-earth and alkaline-earth-like ions with an $ns\ {}^{2}\mathrm{S}_{1/2}$ ground state [5]. While the level structure of such elements offers relatively convenient schemes for laser cooling, state preparation and measurement (SPAM), they lack a large number of accessible internal states and multilevel schemes. In contrast, additional long-lived states in non-alkali-like ions could enable architectures that partition roles across distinct qubits (e.g., processing vs. memory), as proposed in the optical, metastable, and ground-state (omgomg) qubit protocol [1, 41]. While the feasibility of laser cooling of open-$4f$-shell rare-earth ions has been discussed before [18], no real experimental efforts for laser cooling have been demonstrated. Here we present such efforts in $^{169}\text{Tm}^{+}$.
@@
Thulium is a rare-earth lanthanide metal with a single stable isotope, $^{169}\text{Tm}$. Neutral Tm has recently been laser cooled [37, 29] and used for atomic clocks [36, 9], Bose–Einstein condensates [16, 6, 39], and as a quantum-computing platform [26]. $^{169}\text{Tm}^{+}$ in its ground state has a single hole in the $4f$ shell and one electron in the $6s$ shell, generating the $4f^{13}6s\,{}^{3}\mathrm{F}_{4}^{\mathrm{o}}$ state, or $(\nicefrac{{7}}{{2}},\nicefrac{{1}}{{2}})^{\mathrm{o}}_{4}$ in the $J_{1}j$-coupling basis [17]. This configuration yields manifolds of relatively low-lying states with total electronic angular momentum spanning from $J=1$ to $J=9$. Such large-$J$ manifolds are attractive for absorption–emission code schemes that require at least $J\geq\nicefrac{{9}}{{2}}$[13, 2]. At the same time, the nuclear spin $I=\nicefrac{{1}}{{2}}$ ensures that each fine-structure level splits into only two hyperfine-structure (HFS) states, simplifying laser cooling and qubit initialization by reducing the number of frequency tones required for closed population transfer.
@@
Figure 1: E1 transitions and hyperfine splittings in $^{169}\text{Tm}^{+}$. Levels with the same $J$ value are aligned vertically, with $F$ values labeling the hyperfine sublevels. Level energies are indicated in cm-1 and the measured hyperfine splittings are given in GHz. The level spacings are not to scale, and levels irrelevant to this work are omitted. The vacuum wavelength and corresponding frequency are shown next to one of the two strongest hyperfine components ($\Delta F=\Delta J$) of each transition. Gray levels and gray dash-dotted lines indicate levels and expected transitions, respectively, that were searched for but not observed in this work.
@@
For this study, two optical cycling schemes were investigated that have cooling transitions at wavelengths of $450$ nm and $313$ nm. These were selected based on their transition strengths, leakage channels and the accessibility of the required wavelengths with current laser technology. The energy levels and EE1 transitions investigated and relevant for laser cooling and SPAM are depicted in Figure 1. In this diagram, solid arrows indicate transitions observed experimentally, while dashed arrows denote potential decay channels predicted by theoretical calculations but not observed in our spectroscopic measurements. This level diagram is non-exhaustive; for clarity, we omit some intermediate levels and those with $J<3$ and $J>6$. For each wavelength, the measured frequency of one of the two strongest hyperfine components ($\Delta F=\Delta J$) is indicated together with the corresponding inferred level splittings.
Of the spectroscopic data on $^{169}\text{Tm}^{+}$ that has been collected experimentally [23, 22, 42, 15, 11], only three transitions relevant to our proposed cooling schemes have been reported in the literature: $448.25$ nm, $453.07$ nm, and $313.22$ nm, with no data on the transitions corresponding to possible repumping transitions. Guided by theoretical predictions for possible decay channels [32, 33, 28], we measured the transitions’ wavelengths and HFS of the relevant levels.
We employed three different experimental setups to determine the HFS of the levels involved in the proposed cycling schemes. Our most precise measurements were obtained from laser-induced fluorescence on a trapped ion cloud, while preliminary spectra were recorded using a hollow-cathode lamp (HCL) and a small vacuum chamber with an ablation target. The latter two methods can be implemented quickly and cost-effectively compared with an ion-trap apparatus and are therefore well-suited for evaluating the feasibility of working with a new ion species before committing to modifications of an established trap setup. The experimental methods are described in detail in Methods. The following introduction of the proposed laser-cooling cycles is presuming an ion trap environment.
Laser cooling at 450 nm
Figure 2: Population dynamics in $^{169}\text{Tm}^{+}$. (a) Solution of the rate equations for the 450-nm cycle. State labels are given in cm-1. Dashed lines are without repumping lasers, solid lines include $749$, $846$ and $855$-nm repumping lasers. Unknown Einstein $A_{21}$-coefficients were taken from the DREAM database [32, 33, 28]. (b) Experimental scattering rate without the infrared repumping lasers at a He pressure of $4.7\cdot 10^{-6}$ Torr. Blocking the $450$-nm lasers for 22 minutes pauses the optical population transfer. (c) Increase of the steady-state scattering rate with He-buffer-gas pressure in the $313$-nm cycle.
When searching for a closed optical cycle through the ground state at a convenient wavelength, transitions near $450$ nm appear to be an attractive choice. Transitions at $448$ and $453$ nm connect the even-parity 22 308.8222\,308.82-cm-1 state ($4f^{12}5d6s$, $J=4$) to the odd-parity ground and $236.95$-cm-1 states, respectively. According to the DREAM database, the 22 308.8222\,308.82-cm-1 state decays into an odd-parity state at 8 957.478\,957.47-cm-1 through an $E1$ transition at $749$ nm with an Einstein $A$ coefficient of $14.4\cdot 10^{3}$ s-1[32, 33, 28]. Because this state has the same parity as the ground states, $E1$ decays back to the cooling cycle are forbidden. Based on the electron $g$ factor and angular momentum geometry, we estimate a lifetime of $\approx\!90$ ms for this state, which implies the need for repumping lasers.
[…]
We find that adding helium buffer gas greatly increases the observed signal in both 313 and 450-nm cycles. This experimental observation is consistent with the cycles being open, as the buffer gas could be quenching undiscovered, long-lived leakage channels. A second possibility is that the buffer gas simply aids in trapping, leading to a higher ion density. The problem of leakage in the 450-nm cycle is further illustrated by a rate-equation simulation performed with the qspec Python package [27]. As is shown in Figure 2(a), it predicts rapid population loss without repumping. Without the infrared repumping lasers, the population is first driven into the $8957.47$-cm-1 and 20 228.7520\,228.75-cm-1 states on a sub-millisecond timescale. Both states continuously decay back into the ground state, while about $38$ % of the 20 228.7520\,228.75-cm-1 state branches into the long-lived $12457.29$-cm-1 state on a timescale of $20$ s. The $22308.82$-cm-1 state, which is the only state contributing to the fluorescence signal, is not populated significantly after approximately $5$ ms. However, a small amount of less than $1$ % remains until the $12457.29$-cm-1 state is populated. These simulations qualitatively agree with the experimental findings in Figure 2(b), showing the slowly decaying fluorescence signal from this remaining population when the pump lasers are on, but no population loss when the light is blocked. The 20 228.7520\,228.75-cm-1 state being the bottleneck explains why no effect of the $749$-nm laser on the fluorescence signal is observed when only the $12457.29$-cm-1 state is repumped. Consequently, the full optical cylcling scheme is expected to require $749$, $846$ and $855$-nm repumping lasers, reaching steady-state population after about $30$ ms, with $23$ % of the population contributing to the fluorescence signal. Discrepancies between simulation and experiment in the population transfer timescale and observed scattering rate suggest that additional $M1$ decay channels, not listed in the DREAM database, might need to be considered. Figure 2(c) shows the increased fluorescence signal when using He-buffer gas, here measured with the $313$-nm transition; hinting at quenching of the metastable states.
Laser cooling at 313 nm
The $313$-nm line is the strongest transition we measured, with an Einstein coefficient of $A_{21}=2\pi\times 16.9$ MHz [17]. It connects the ground state to an even-parity $J=5$ state with a $4f^{12}5d6s$ electronic configuration at 31 926.8231\,926.82 cm-1. Its relatively large linewidth, together with its short transition wavelength, offers a high Doppler cooling rate.
Based on $E1$ selection rules, some population from the $313$-nm cooling cycle is expected to decay into the 22 457.5122\,457.51-cm-1 $(\nicefrac{{7}}{{2}},\nicefrac{{3}}{{2}})_{4}^{\mathrm{o}}$] and 23 524.0923\,524.09-cm-1 $(\nicefrac{{7}}{{2}},\nicefrac{{5}}{{2}})_{4}^{\mathrm{o}}$] levels. We applied light at $854.69$, $904.43$, $1056.04$, and $1190.09$ nm to repump population through different excited states within the $313$-nm cycle (see Figure 1), but observed no measurable change in fluorescence. This indicates that either the corresponding decay channels are weaker than the $854.84$-nm channel, or the lifetimes of these intermediate states are too short to cause significant population trapping.
Another potential metastable $(6,1)_{5}$ state suggested by the DREAM database [32, 33, 28] is located at 16 567.4716\,567.47 cm-1. Based on its energy and selection rules, this level should decay to the ground state via an EE1 transition at approximately $604$ nm. However, driving the 1 297.741\,297.74-nm transition connecting this state to the 24 273.2024\,273.20-cm-1 level produced no observable change in the fluorescence signal.
Laser cooling on other transitions
A $370$-nm transition connecting the ground state to a $J=4$ state at 27 254.4227\,254.42 cm-1 (not shown in Figure 1) is another candidate for laser cooling. While this was the first transition we observed in the HCL, it is estimated to be $5$ times weaker than the $313$-nm transition. Furthermore, it shares most of the decay pathways to higher-$J$ states and introduces new decay channels toward $J\leq 3$ metastable states. The $313$-nm line is therefore considered a more feasible option.
While this work has focused on transitions from the ground state to $J=4$ and $J=5$ manifolds, other configurations are possible. For example, a recent UV spectroscopy study [11] identified several promising transitions originating from the 8 957.478\,957.47-cm-1 state, including transitions to $J=3$ states near $363$ and $331$ nm with high branching factors of $0.995$ and $0.971$, respectively. Although lasers at these specific wavelengths were not available for the present work, these transitions represent a compelling avenue for future investigation.
Sympathetic cooling with $^{138}\text{Ba}^{+}$
Our ion trap can reliably trap and laser cool large clouds of $^{138}\text{Ba}^{+}$ ions. By repeatedly switching the trap potential between high and low values, ions can be selectively removed to produce a laser-cooled few-ion crystal or a single ion. In principle, simultaneously co-trapping $\text{Ba}^{+}$ and $\text{Tm}^{+}$ should allow for the sympathetic cooling of $\text{Tm}^{+}$ ions during $\text{Ba}^{+}$ laser cooling. While we have demonstrated co-trapping of large $\text{Ba}^{+}$ and $\text{Tm}^{+}$ clouds via trap-door loading, attempts to crystallize the two species simultaneously, so far, resulted in the loss of all $\text{Tm}^{+}$ ions. We attribute this behavior to the lower mass of the $\text{Ba}^{+}$ ions, which experience a larger ponderomotive force. Consequently, they tend to push the $\text{Tm}^{+}$ ions outwards during laser cooling rather than uniformly mix with or integrate the $\text{Tm}^{+}$ ions into the ion crystal. Because of this spatial separation, we tend to lose $\text{Tm}^{+}$ ions faster when toggling the trap potential. A potential solution is to utilize ions of a more similar mass, such as 171Yb+ for sympathetic cooling; however, this capability is currently not available in our experimental setup. An alternative approach is to first trap and crystallize $\text{Ba}^{+}$ and then introduce $\text{Tm}^{+}$ without lowering the trap potential, either by resonant photo-ionization of neutral Tm or by capturing $\text{Tm}^{+}$ through collisions with the trapped $\text{Ba}^{+}$ ions.
[…]
Results
Hyperfine spectroscopy
@@ Results
A selection of fluorescence spectra that resulted from our measurement campaign is depicted in Figure 3. From these spectra, the hyperfine-structure (HFS) $A$ constants of the excited states relevant to the proposed cooling schemes were determined. The extracted values are listed in Table 1. Each spectrum was fitted with a sum-of-Gaussians model, using the HFS $A$ parameters of the lower and upper states as fit variables. Due to the large total electronic angular momentum, transitions where $\Delta F\neq\Delta J$ are strongly suppressed. As a result, in several cases, only two spectral features were available to extract a single unknown $A$ constant, with the other fixed to a known reference value.
Figure 3: Hyperfine-structure (HFS) spectra of $^{169}\text{Tm}^{+}$. (a,b) were recorded using absorption spectroscopy with a hollow-cathode lamp, and (c–i) using fluorescence spectroscopy with an ionic cloud in a linear quadrupole trap. (c–f) A fresh ion cloud was produced for each data point; the laser light was generated using external-cavity diode lasers (ECDLs) and the laser frequencies were set to fixed values for each point. (g–i) Each spectrum was recorded with a single cloud of ions using the continuous-scan function of the optical parametric oscillator (OPO) laser.
The specified uncertainties were determined from the least-square-fit uncertainties, adjusted for a reduced $\chi^{2}$ of one, as well as the statistics of repeated measurements and the frequency readout. A statistical $5$ MHz frequency-readout uncertainty was added in quadrature (geometric sum) to the uncertainties of the determined resonance peak positions of spectra generated with the ECDLs, before calculating the weighted average over all measurements. The maximum between the standard errors of the unweighted and weighted average was then taken as the total statistical uncertainty. In cases where the $A$ constant was predominantly extracted from continuous scans of our OPO laser, a $2$ MHz systematic frequency-calibration uncertainty was added linearly to the statistical uncertainty. This systematic uncertainty is the dominant contribution to our error budget. Other potential sources, such as light shifts, photon recoil shifts, and line-shape asymmetries, are negligible at the current level of precision. It should be noted that our uncertainty estimates are conservative, as performing metrological-grade high-precision measurements remains beyond the scope of this work.
@@
Table 1: Hyperfine-structure (HFS) constants in $^{169}\text{Tm}^{+}$. The doublet splittings are readily calculated using $(J+\nicefrac{{1}}{{2}})A$. The HFS constants of the ground, $236.95$, and 31 926.8231\,926.82-cm-1 states were kindly provided by Ref. [4]. The precise value of the 12 457.2912\,457.29-cm-1 state was determined using microwave spectroscopy.